OpenStudy (anonymous):

A basket contains the following pieces of fruit: 3 apples, 2 oranges, 2 bananas, 2 pears, and 5 peaches. Jonas picks a fruit at random and does not replace it. Then Beth picks a fruit at random. What is the probability that Jonas gets a peach and Beth gets an apple?

OpenStudy (nathan917):

I will assume that the order of Jonas picking, then Beth is important. so prob(as stated) = (5/11)(4/10) = 2/11